OpenStudy (mathmath333):

A dog after travelling 50 km meets a swami who counsels him to go slower.He then proceeds at 3/4th of his formal speed and arrives 35 min late.Had the meeting occured 24 km further the dog would have reached the destination 25 mins late.The speed of the dog is? a.)48 km/hr b.)36 km/hr c.)54 km/hr d.)58 km/hr

OpenStudy (rational):

what meeting the dog is attending ?

OpenStudy (mathmath333):

he is attending a secret meeting with a swami

OpenStudy (rational):

Also is it an indian dog ?

OpenStudy (mathmath333):

yep

OpenStudy (rational):

lmao

OpenStudy (rational):

lets say the original speed of dog is \(s\)

OpenStudy (rational):

say the distance is \(d\) : \[\dfrac{50}{s} + \dfrac{d-50}{\frac{3}{4}s} = \dfrac{d}{s}+35\tag{1}\]

OpenStudy (mathmath333):

 is dog travelling back 50 km at \(3s/4\) speed

OpenStudy (rational):

Okay looks i misinterpreted the problem let me try again :)

rishavraj (rishavraj):

@mathmath333 hey i got the answer..... and plz expalin where its written tht he is returning back

rishavraj (rishavraj):

*explain

OpenStudy (mathmath333):

this page has a solution , which i didnt get http://www.piit.testfunda.com/Answers/ViewQuestion.aspx?QID=0fac7f0e-ea91-4869-a85f-d73caca87114

rishavraj (rishavraj):

@mathstudent55 mine answer is correct

OpenStudy (mathmath333):

oh sry ,it is not written that he is returning back

OpenStudy (rational):

hmm i don't like interpreting other's solutions much... let me try once again

rishavraj (rishavraj):

see let the whole distance be x km.....and his speed be "a" km/hr so if he travels whole distance at tht speed he is gonna take time \[T _{1} = x/a\]

OpenStudy (mathmath333):

i am assuming now that he is going forward and not returning back

rishavraj (rishavraj):

@mathmath333 u need to frame few equations using the data u r provided.......:)

OpenStudy (mathmath333):

yep i know, i tried and failed.

OpenStudy (mathmath333):

because i cant understand the question correctly

rishavraj (rishavraj):

see the time he would take if he travelled at "a" km/hr is T = x/a now see he travels 50 km with tht speed...so time he takes to travel tht distance T1 = 50/a hr now the distance left i.e (x - 50) he travells at (3/4)a km/hr....si time taken T2 - (x - 50)4/3a

rishavraj (rishavraj):

question says T1 + T2 = T + 35/60 u need to change 35 mins into hour

rishavraj (rishavraj):

try to frame another equation urself........@mathmath333 :)))

OpenStudy (mathmath333):

u mean \(\Large T_2 = \frac{(x - 50)4}{3a}\) ?

rishavraj (rishavraj):

it says "Math Processing error" lol...;P

OpenStudy (mathmath333):

reload the page

rishavraj (rishavraj):

nah brother its failing.......hold on i m uploading a pic ogf the solution....:))

OpenStudy (mathmath333):

and second equation is this ? \(\large \color{black}{\begin{align} \dfrac{74}{a}=\dfrac{4(x-74)}{3a}+\dfrac{25}{60}\hspace{.33em}\\~\\ \end{align}}\)

rishavraj (rishavraj):

the first equation is \[(x - 50)4/3a + 50/a = x/a + 35/60\]

rishavraj (rishavraj):

give another try for the second equation see this he travels 74 km with speed "a" km/hr

rishavraj (rishavraj):

its not tht....see T3 = 74/a and the rest distance i.e (x - 74) he travels at (3/4)a km/hr so T4 = (x - 74)4/3a

OpenStudy (mathmath333):

i have attached by second equation in pic

rishavraj (rishavraj):

T3 + T4 = x/a + 25/60

rishavraj (rishavraj):

@mathmath333 ??????

OpenStudy (mathmath333):

yes

rishavraj (rishavraj):

solve the equations u will get the answer.....:))

OpenStudy (mathmath333):

what answer did u get from there

OpenStudy (mathmath333):

option a . 48 ?

rishavraj (rishavraj):

48 km/hr.....:))

rishavraj (rishavraj):

wbu????

OpenStudy (mathmath333):

can u write again ur two equations in one line .

rishavraj (rishavraj):

first equation \[\frac{ (x - 50)4 }{ 3a } + \frac{ 50 }{ a } = \frac{ x }{ a } + \frac{ 35 }{ 60 }\]

OpenStudy (mathmath333):

ok

rishavraj (rishavraj):

second \[\frac{ (x - 74)4 }{ 3a } + \frac{ 74 }{ a } = \frac{ x }{ a } + \frac{ 25 }{ 60 }\]

OpenStudy (mathmath333):

ok thanks i will ponder this equations now

rishavraj (rishavraj):

all the best